OpenStudy (madeleine2):

Write the point-slope form of the line that passes through (1, -5) and is parallel to a line with a slope of 1. Include all of your work please.

OpenStudy (idkanything):

@3mar

OpenStudy (madeleine2):

i think the answer is y=x-6 but i used an algebra calculation and i have no idea how to get to the answer

OpenStudy (idkanything):

3mar is smart and probably knows

OpenStudy (3mar):

Well, I am here.

OpenStudy (3mar):

Thanks for your confidence, @Idkanything! Honor is mine!

OpenStudy (idkanything):

good we needed u

OpenStudy (idkanything):

ha ha

OpenStudy (madeleine2):

can I get some help i dont know how to do this question

OpenStudy (3mar):

Won't be late for any help.

OpenStudy (idkanything):

3mar do u know?

OpenStudy (3mar):

Yes.

OpenStudy (3mar):

The point-slope form of the line that passes through a point (x1,y1) and has the slope m is \[y-y_1=m(x-x_1)\]

OpenStudy (madeleine2):

yes i know that the formula is y−y1=m(x−x1) but i dont know how to go from there once i plug in the numbers which i think is y--5=1(x-1)but what do you do after this

OpenStudy (3mar):

Can you get the slope of the line which is parallel to the line with the slope 1?

OpenStudy (madeleine2):

i dont know how to do that

OpenStudy (3mar):

 for any two or more parallel lines, their slopes are the same.

OpenStudy (madeleine2):

i really dont get this can you just tell me the steps to get to the final equation

OpenStudy (madeleine2):

@3mar

OpenStudy (3mar):

If you get the slope from the parallel line, it would be also 1 because the y are parallel and they have the same slope

OpenStudy (3mar):

then plug in these two values of slope and point in the form I typed for you

OpenStudy (3mar):

\[y-y_1=m(x-x_1)\] \[y-(-5)=(1)(x-1)\] \[y+5=x-1\] \[y=x-6\]

OpenStudy (3mar):

This also could help you. https://www.desmos.com/calculator/j9xfuztxcf

OpenStudy (3mar):

@madeleine2

OpenStudy (madeleine2):

thank you so so much

OpenStudy (madeleine2):

i tried to do that with this equation but it isnt working Write the point-slope form of the line that passes through (5, 5) and is parallel to a line with a slope of 1/4. Include all of your work in your final answer.

OpenStudy (3mar):

You are so welcome!

OpenStudy (3mar):

"i tried to do that with this equation but it isnt working" why?

OpenStudy (3mar):

@madeleine2 please I am busy and you let me wait a lot!